# Evaluating Perspectival Biases in Cross-Modal Retrieval

### v1

@@ 4 Experimental Setup
To answer our research questions, we conducted two main experiments. For RQ1, we performed image-to-text retrieval on the Crossmodal-3600 dataset [23]. The dataset offers a balanced multilingual text pool comprising native captions in 36 languages, making it suitable for auditing cross-lingual behavior in image–text retrieval, without implying any particular pattern of disparities. We evaluate models using Accuracy@5, NDCG@10, LBKL@10, and our proposed DLBKL@10. For RQ2, we performed text-to-image retrieval on our newly created XCM benchmark, evaluating models using our proposed SP score.
For both RQ1 and RQ2, we selected a representative suite of models spanning three distinct architectural paradigms, as shown in Table 1:
  • Vision-Language Contrastive Models: These are foundational models trained primarily on English data. We include the original CLIP-L/14 as a powerful baseline, and Chinese-CLIP-L/14 to observe the effect of monolingual fine-tuning on a non-English corpus.
  • Cross-lingual Alignment Models: These models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space. We evaluate two variants of m-CLIP, which use XLM-RoBERTa as the text encoder (XLM-R-L/14 and XLM-R-B/16plus).
Full model identifiers are available in Appendix H.
@@ 5.1 Image-to-Text Evaluation (RQ1)
All models exhibit some degree of linguistic prevalence bias. For most MLLM-based models (Jina, ColQwen, GME), the DLBKL score is higher than the LBKL score, as shown in Table 1.This confirms that bias is more pronounced at the top of the ranked list, as these models tend to rank results from medium-to-high resource languages. Results for additional ranks and an example of retrieval result can be found in Appendix D.
@@
ColQwen2.5-3b-M 0.894 0.792 0.817 0.605
ColQwen2.5-7b-M 0.926 0.821 0.849 0.665
ColQwen2.5-v0.2 0.754 3.834 3.867 0.481
GME-Qwen2-2B 0.967 3.121 3.174 0.717
GME-Qwen2-7B 0.979 1.371 1.420 0.770
Jina-E-v4 0.972 0.915 0.951 0.775
Table 1: Image-to-text retrieval on Crossmodal-3600. Bias is measured by LBKL and DLBKL. Explicit alignment models (XLM-R) show substantially lower bias.
This phenomenon is visualized in Figure 6, which shows a clear dominance of high-resource languages in the top ranks. This further illustrates the overall disparity in retrieval frequency between language resource tiers as shown in Figure 7.
Figure 6: Distribution of language groups across retrieval ranks. High-resource languages (blue) dominate the top ranks, a bias captured by DLBKL.
Figure 7: Histogram of retrieved language frequencies. MLLM-based models disproportionately retrieve texts from medium-high resource languages (blue) over low-resource ones (orange).
@@ 5.2 Text-to-Image Evaluation (RQ2)
Using the proposed XCM benchmark, we evaluated the association bias of several multimodal retrievers, ranging from CLIP to more recent models. In this evaluation, the semantic win rate ($M_{\text{sem}}$) serves as a proxy for raw performance, while the SP score quantifies cultural bias. We observe that the baseline CLIP and Chinese-CLIP models exhibit a significant cultural bias, often preferring a culturally associated but semantically incorrect image, as shown in Table 2.
@@
LLM-Based Retrieval Embedders
@@
Our culture-specific analysis reveals that this self-preference is a symptom of missing linguistic knowledge, as shown in Figure 8. The CLIP-L/14 model, lacking a robust understanding of non-Latin scripts, defaults to matching cultural origin as a retrieval heuristic. Training on a large Chinese dataset (Chinese-CLIP) partially addresses this, improving performance for both Chinese and Japanese queries due to the shared logographic Kanji characters. However, this is a shallow fix that fails to generalize to other non-Latin scripts. In contrast, the text-aligned model (XLM-R-L/14) performs well across most languages, with a notable exception for queries in Yoruba (Nigeria). This challenge with low-resource languages persists even in more advanced architectures. For instance, MLLM-based models (Jina-E-v4) employ a LLM as their text encoder, leveraging its pre-training on web-scale multilingual data for a robust understanding of diverse languages. For the vision component, a Vision-Language Model (VLM) is used as the image encoder to improve contextual awareness. However, performance drops for low-resource languages.
@@
Figure 8: Association bias evaluation across four models reveals the limitations of monolingual training. The baseline CLIP (a) shows significant cultural bias, which is exacerbated by region-specific fine-tuning as seen in Chinese-CLIP (b). In contrast, cross-lingual models like XLM-R (c) and particularly Jina-E-v4 (d) prove far more effective at mitigating this bias and maintaining high semantic relevance across diverse countries.
@@
Figure 9: UMAP projection of native concept embeddings across four models: (a) CLIP-L/14 (non-Latin language separation), (b) Chinese-CLIP-L/14 (language family clustering), (c) XLM-R-L/14 (dense single-cluster unification), and (d) Jina-E-v4 (unified but dispersed cluster).
@@
This behavior is clearly visualized in the UMAP [24] projections of the text embeddings as shown in Figure 9. The baseline CLIP-L/14 model exhibits a fractured embedding space, with non-Latin languages forming distinct clusters far from the main Latin-script cluster. This demonstrates a lack of shared semantic understanding. In the Chinese-CLIP model, the Chinese and Japanese embeddings shift closer to the Latin cluster, reflecting the targeted training, but other non-Latin languages remain isolated. In contrast, the explicit alignment model, XLM-R-L/14, successfully unifies the embedding space into a single, language-agnostic cluster, demonstrating a truly shared semantic representation across scripts. The only notable outlier is Yoruba, which was not part of this specific model’s alignment training. The MLLM model, Jina-E-v4, exhibits a similar but distinct pattern: it also forms a single, unified cluster, but the embeddings are more widely dispersed. This suggests a more flexible alignment that may capture finer semantic nuances between languages.
@@
To validate these visual findings numerically, we calculated the silhouette score [25] for each language’s text embeddings. This analysis revealed a strong Pearson correlation (0.68) between a language’s silhouette score and its measured SP score as shown in Appendix I. This quantitatively reinforces that poor semantic understanding in the text encoder (as visualized by the disparate UMAP clusters) is a key driver of higher cultural association bias.
@@ 6 Discussion
Our evaluation framework distinguishes between two perspectival biases: prevalence bias, driven by data imbalance, and association bias, arising from learned cultural correlations. Our findings show these are distinct challenges.
Explicit cross-lingual alignment, used by the XLM-R models, is a highly effective strategy, achieving the lowest scores for both prevalence bias (DLBKL) and association bias (SP) by directly enforcing a shared semantic space. While modern MLLMs like Jina-E-v4 also perform well against association bias, the persistence of these issues across all models points to a deeper, unresolved problem: the entanglement of semantic concepts with linguistic and cultural artifacts in the model’s embedding space.
The path forward, therefore, requires a fundamental shift in training strategy. Future work must prioritize training objectives that actively enforce non-association by creating a truly global semantic space. This means designing models to map a semantic query, regardless of its language or cultural origin, to all conceptually relevant images, irrespective of their geographical context. For example, the new method must include a curation process to avoid cross-cultural false negative images being pushed away from their corresponding queries, and utilize data augmentation to help ensure language-agnostic property.
@@ 7 Conclusion
In this work, we introduce a framework that distinguishes between two forms of perspectival bias in multimodal retrieval, prevalence bias and association bias, reflecting distinct ways in which a model’s prior shapes its behavior. This conceptual framing is operationalized through our proposed metrics and datasets: DLBKL, which measures rank-aware language prevalence bias, and XCM, which quantifies association bias through cross-cultural image retrieval. Together, these tools enable systematic evaluation of how multimodal large language models inherit and express perspectival biases across languages and cultures.
In image-to-text retrieval, prevalence bias arises when a model favors texts from high-resource languages. This problem is addressed by anchoring other languages to the prevalent ones through cross-lingual alignment. In text-to-image retrieval, association bias arises when a model favors images that are culturally associated with the query language rather than semantically faithful to their content. Such bias cannot be resolved through traditional cross-lingual alignment or by merely exposing the model to a wider range of cultural content during training. Ultimately, our findings call for a more principled approach: one that directly mitigates localized spurious association as a core design principle for models that are not only multilingual but also perform consistently across languages and cultures.
@@ 8 Limitations
Our work has several limitations. First, our DLBKL metric measures fairness via distributional parity, not semantic correctness. It therefore cannot distinguish between retrieving irrelevant documents and over-representing a language with relevant ones. Second, the XCM benchmark simplifies culture by using country as a proxy, a necessary choice for tractability that does not capture transnational or sub-national cultures. The benchmark’s coarse-grained semantics (e.g., "food") and lack of accounting for polysemy also limit its representation of real-world query complexity.
Acknowledgments
This research has received funding support from the National Science, Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation [grant number B13F680099] and Chulalongkorn University Artificial Intelligence Institute.
Appendix A Culturally Relevant Concept Identification
To identify image concepts unique to each country, we first employed the Gemini2 as a tool for generating culturally relevant suggestions prior to data collection. The prompt used in this process is shown in Figure 11 to identify all labels associated with an image. Figure 10.
list 100 concepts that unique and vary in these country including China, India, Japan, saudi arabia, France, German, Brazil,
Kenya, Thailand, USA
like this
{"food":{"China":"Mala Xiang Guo", ..., "Thailand":"Padthai", "USA":"hamburger"}, "costume":{"China":"Hanfu", ...,
"Thailand":"Sabai", "USA":"cowboy"}}
Appendix B Multi-Concept Detection in Cultural Images
Please classify the following image by assigning them to one or more of the following cultural
categories:
   {category}.
**Comprehensive Output Format (JSON):**
output as JSON for example:
  "<INDEX>": ["<CATEGORY>", "<CATEGORY>", ...],
in the categories please order by priority (high to low).
Appendix C Language Resources
To estimate language resource availability, we utilized the Distribution of Languages from the Common Crawl dataset (CC-MAIN-2025-18) [26] as an approximation. Table 3 and 4 presents the resulting language composition for RQ1 and RQ2, consequently.
Language Type Language ID Distribution (%)
@@ Appendix C Language Resources
Name Full Name Language Resources
USA United States of America English 43.950
UK United Kingdom
AUS Australia
GER Germany German 5.569
CHN China Chinese 4.878
JPN Japan Japanese 4.915
ESP Spain Spanish 4.542
ARG Argentina
FRA France French 4.327
PRT Portugal Portuguese 2.337
BRA Brazil
SAU Saudi Arabia Arabic 0.672
THA Thailand Thai 0.427
IND India Hindi 0.200
KEN Kenya Swahili 0.010
NGA Nigeria Yoruba 0.001
@@
Appendix D Example of Result from Image to Text Retrieval
To further elaborate the result of research question 1, we provide the example of retrieval from image to text from CLIP and M-CLIP in Figure 12. We also provide result of LBKL and DLBKL score at other rank in Table 5.
Figure 12: An Example of Result from Image to Text Retrieval
Model @5 @25 @50 @99
[…]

### v3

@@ 4 Experimental Setup
To answer our research questions, we conducted three main studies. For RQ1, we performed image-to-text retrieval on the Crossmodal-3600 dataset [ThapliyalCrossmodal2022]. The dataset offers a parallel multilingual text pool comprising native captions in 36 languages, making it suitable for auditing cross-lingual behavior in image–text retrieval, without implying any particular pattern of disparities. With this dataset, we compute Language-wise DLBKL by defining the prior $P(x)$ as a uniform distribution (see Appendix E)". We evaluate models using Accuracy@5, NDCG@10, LBKL@10, and our proposed DLBKL@10. For RQ2, we performed text-to-image retrieval on our newly created 3XCM benchmark, evaluating models using our proposed SP score. For RQ3, we extend text-to-image retrieval of RQ2 by using the extended version of the 3XCM benchmark with six candidates and analyze the similarity drift.
We selected a representative suite of models spanning three distinct architectural paradigms: Vision-Language Contrastive Models: foundational models trained with a contrastive objective mainly on English data, Cross-lingual Alignment Models: models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space for multilingual capability, and MLLM-Based Retrieval Embedders: utilize MLLM as the backbone and finetune for a contrastive objective for retrieval. We refer readers to Appendix C for detailed architectural descriptions and specific model configurations.
@@ 5.1 Image-to-Text Evaluation (RQ1)
All models exhibit some degree of linguistic prevalence bias, as shown in Table 1. This bias is most pronounced in the top-ranked results, as models tend to prioritize high-resource languages over others in the early ranks. Appendix G provides visualizations revealing the dominance of high-resource languages in the top ranks and the overall disparity in retrieval frequency between language resource tiers. Results for additional ranks and an example of a retrieval result can be found in Appendix F.
@@
ColQwen2.5-3b-M 0.894 13.654 13.667 0.605
ColQwen2.5-7b-M 0.926 13.439 13.454 0.665
ColQwen2.5-v0.2 0.754 14.228 14.241 0.481
GME-Qwen2-2B 0.967 13.627 13.644 0.717
GME-Qwen2-7B 0.979 13.378 13.397 0.770
Jina-E-v4 0.972 13.008 13.025 0.775
Table 1: Image-to-text retrieval on Crossmodal-3600. Bias is measured by LBKL and DLBKL, calculated in a language-wise manner. Explicit alignment models (XLM-R) show substantially lower bias.
@@ 5.2 Text-to-Image Evaluation (RQ2)
Using the proposed 3XCM benchmark, we evaluated the association bias of several multimodal retrievers, ranging from CLIP to more recent models. In this evaluation, the semantic win rate ($M_{\text{cr}}$) serves as a proxy for raw performance, while the SP score quantifies cultural bias. We observe that the baseline CLIP and CN-CLIP models exhibit a significant cultural bias, often preferring a culturally associated but semantically incorrect image, as shown in Table 2.
@@
MLLM-Based Retrieval Embedders
@@
Our culture-specific analysis reveals that this self-preference is a symptom of missing linguistic knowledge, as shown in Figure 5. The CLIP-L/14 model, lacking a robust understanding of non-Latin scripts, defaults to matching cultural origin as a retrieval heuristic. Training on a large Chinese dataset (CN-CLIP) partially addresses this, improving performance for both Chinese and Japanese queries due to the shared logographic Kanji characters. However, this is a shallow fix that fails to generalize to other non-Latin scripts. In contrast, the text-aligned model (XLM-R-L/14) performs well across most languages, with a notable exception for queries in Yoruba (Nigeria). This challenge with low-resource languages persists even in more advanced architectures. For instance, MLLM-based models employ an LLM as their text encoder, leveraging its pre-training on web-scale multilingual data for a robust understanding of diverse languages. For the vision component, a Vision-Language Model (VLM) is used as the image encoder to improve contextual awareness. However, performance drops for low-resource languages.
@@
Figure 5: Association bias evaluation across four models reveals the limitations of monolingual training. The baseline CLIP shows significant cultural bias, which is exacerbated by region-specific fine-tuning as seen in CN-CLIP. In contrast, cross-lingual models like XLM-R and particularly Jina-E-v4 prove far more effective at overriding this bias and maintaining high semantic relevance across diverse countries.
@@
Figure 6: UMAP projection of native concept embeddings across four models: (a) CLIP-L/14 (non-Latin language separation), (b) CN-CLIP-L/14 (language family clustering), (c) XLM-R-L/14 (dense single-cluster unification), and (d) Jina-E-v4 (unified but dispersed cluster).
@@
This behavior is clearly visualized in the UMAP [McInnes2018] projections of the text embeddings as shown in Figure 6. The baseline CLIP-L/14 model exhibits a fractured embedding space, with non-Latin languages forming distinct clusters far from the main Latin-script cluster. This demonstrates a lack of shared semantic understanding. In the CN-CLIP model, the Chinese and Japanese embeddings shift closer to the Latin cluster, reflecting the targeted training, but other non-Latin languages remain isolated. In contrast, the explicit alignment model, XLM-R-L/14, successfully unifies the embedding space into a single, language-agnostic cluster, demonstrating a truly shared semantic representation across scripts. The only notable outlier is Yoruba, which was not part of this specific model’s alignment training. The MLLM model, Jina-E-v4, exhibits a similar but distinct pattern: it also forms a single, unified cluster, but the embeddings are more widely dispersed. This suggests a more flexible alignment that may capture finer semantic nuances between languages.
@@
To validate these visual findings numerically, we calculated the silhouette score [shahapure2020cluster] for each language’s text embeddings. This analysis revealed a strong Pearson correlation (0.68) between a language’s silhouette score and its measured SP score as shown in Appendix N. This quantitatively reinforces that poor semantic understanding in the text encoder (as visualized by the disparate UMAP clusters) is a key driver of higher cultural association bias.
@@
Model behavior remains consistent regardless of prompt phrasing, such as verbose descriptions (e.g., “a picture of a train for my homework”) or culture-agnostic instructions (e.g., “focus on semantics instead of textual language”) (Appendices R.1 and S).
5.3 Cultural Descriptor Text-to-Image Evaluation (RQ3)
To investigate whether explicit instruction can override implicit association bias, we evaluate the models on the Extended 3XCM benchmark using queries with explicit Cultural Descriptors (CD) (e.g., “Train of Japan” in Thai).
Adherence to Cultural Descriptors.
Quantitative results demonstrate that most models are responsive to explicit cultural cues. As shown in Table 3, the correct answer (matching both object and cultural descriptor) is the most frequently selected candidate across all high-performing models (XLM series, MLLM-based). This suggests that explicit cultural descriptors (e.g., the word “Japan”) exert a stronger influence on the retrieval ranking than the implicit bias of the query language (e.g., Thai script), provided the model has sufficient multilingual capacity.
Figure 7: Comparative association bias across CLIP-L/14, CN-CLIP-L/14, Jina-embeddings-v4, and XLM-R-L/14 models in 16 countries. Stacked bars show win percentages for six categories, ranging from Correct to Total Irrelevance, highlighting performance variance across diverse cultural contexts.
Model $M_{\text{cr}}$ $M_{\text{orlb}}$ $M_{\text{or}}$ $M_{\text{cdr}}$ $M_{\text{lb}}$ $M_{\text{ti}}$
Vision-Language Contrastive Models
CLIP-L/14 43.05 29.83 5.30 8.54 12.56 0.72
CN-CLIP-L/14 41.97 24.54 9.81 10.04 11.21 2.42
Cross-lingual Alignment Models
XLM-R-L/14 77.83 7.21 6.85 7.55 0.26 0.30
XLM-R-B/16plus 72.98 9.88 9.70 6.57 0.40 0.48
MLLM-Based Retrieval Models
GME-Qwen2-2B 60.01 18.38 11.49 8.05 1.56 0.50
GME-Qwen2-7B 59.75 20.84 10.79 6.81 1.38 0.43
ColQwen2.5-v0.2 54.77 20.15 12.57 9.56 2.13 0.82
ColQwen2.5-3B-M 58.73 20.71 9.26 9.37 1.31 0.62
ColQwen2.5-7B-M 63.19 17.70 7.82 9.80 1.11 0.38
Jina-E-v4 57.76 19.59 13.62 7.14 1.20 0.68
Table 3: Association bias results on the 3XCM benchmark (6 categories). $M_{\text{cr}}$, $M_{\text{orlb}}$, $M_{\text{or}}$, $M_{\text{cdr}}$, $M_{\text{lb}}$, and $M_{\text{ti}}$ represent Correct, Object Relevant with Language Bias, Object Relevant, Cultural Descriptor Relevant, Language-biased, and Totally Irrelevant, respectively.
Failure Mode Analysis.
Let us examine two failure modes presented by the error distribution in Figure 7 First, when the model successfully retrieves the correct semantic object but fails to adhere to the explicit cultural descriptor (i.e., "Semantic Success, Cultural Failure"), we observe that the selection rate of the Object Relevant Language-biased category (which matches the object and the query language’s culture, represented by purple color) is higher than that of the Object Relevant category (which matches the object with an unrelated culture, represented by blue color). This difference is significant ($p$-value < 0.05, Chi-squared test) for most models except the XLM family (see Appendix P.2). The trend is most pronounced in non-Latin and low-resource languages, suggesting that when the explicit cultural constraint is violated, the model falls back on implicit language-driven associations.
Second, when the model fails to retrieve the correct semantic object (i.e., "Semantic Failure"), the model is more likely to select the Cultural Descriptor Relevant category (which matches the explicit cultural descriptor only, represented by yellow color) than the Language-biased category (which matches the query language’s culture only, represented by pink color). This difference is significant ($p$-value < 0.05, Chi-squared test) for most models except the Vision-Language Contrastive group (CLIP and CN-CLIP), which lack robust non-Latin script understanding (see Appendix P.2). This indicates that when the model fails to ground the visual object, it prioritizes the explicit cultural token (e.g., the country name) over the implicit association of the script. (See Appendices P and R.2).
Similarity Drift Analysis.
The Similarity Drift ($\Delta\text{sim}$) metric measures the change in cosine similarity for a target image when a cultural descriptor is added to the query. On average, adding a cultural descriptor successfully shifts the similarity distribution in the positive direction (toward the Correct category) across most models and languages. This confirms that the models generally recognize and attend to the explicit instruction. However, a notable exceptions emerge. For the baseline CLIP model, queries in non-Latin scripts exhibit near-zero drift, indicating that the text encoder fails to process the cultural descriptor in these scripts, rendering the explicit instruction ineffective (see Appendix Q.1).
@@ 6 Discussion
Perspectival Biases and the Power of Explicit Alignment. Our investigation establishes that perspectival bias manifests in two correlated forms: prevalence bias (RQ1) from distributional imbalances in training data, favoring high-resource languages, and association bias (RQ2) from spurious correlations between language scripts and cultural visual features. While both arise from training data and correlate empirically (high DLBKL models tend toward high SP), explicit cross-lingual alignment effectively reduces both, as seen in XLM-R models’ low bias scores and high accuracy. This alignment not only mitigates prevalence and association biases but also enhances the model’s ability to follow explicit cultural descriptors (CDs).
Guidance via Cultural Descriptors (CDs) and the Tugging Effect. CDs provide guidance to steer culture. The positive Similarity Drift ($\Delta\text{sim}$) in RQ3 shows models can adhere to cultural descriptors, steering retrieval toward target cultures in most cases despite the tugging effect of query-language associations. However, failure mode analysis reveals the tugging effect: when matching the object but not the CD, models revert to culture-language association over neutral alternatives. This indicates association bias as a latent prior suppressed but not eliminated by guidance. Modern MLLM-based models still exhibit issues with both biases and inconsistent compliance with CD guidance, especially for low-resource or non-Latin languages, where drifts are erratic or near-zero without robust multilingual support.
@@ 7 Conclusion
We distinguish prevalence bias, a distributional skew, from association bias, a semantic entanglement between language and culture. Even with explicit cultural descriptors, models exhibit a systematic “tugging effect” to language-culture association bias when semantic grounding succeeds but cultural disambiguation fails. These findings demonstrate that achieving cross-cultural equity requires more than naive scaling. Future work must prioritize training paradigms that actively de-correlate language from cultural features, enforce cultural constraints for steerability, or explicit cross-lingual alignment. Only by compiling these can we achieve a retrieval system that truly operates across global contexts with cultural controllability.
@@ 8 Limitations
Our work has several limitations. First, our DLBKL metric measures fairness via distributional parity, not semantic correctness. It therefore cannot distinguish between retrieving irrelevant documents and over-representing a language with relevant ones. Second, the 3XCM benchmark simplifies culture by using country as a proxy, a necessary choice for tractability that does not capture transnational or sub-national cultures. The benchmark’s universal basic-level semantics (e.g., "food") and lack of accounting for polysemy also limit its representation of real-world query complexity, as its usage is intended to be used as a challenging, controlled environment used to uncover biases.
Appendix A Comparison with Other Datasets
We compare 3XCM with existing VQA benchmarks in Table LABEL:tab:benchmark_comparison. The 3XCM dataset was developed to address the lack of cross-cultural parallelism in current resources.
Feature MaRVL [liu-etal-2021-visually] CulturalVQA [nayak-etal-2024-benchmarking] CVQA [romero2024cvqa]
Primary Task Visual Reasoning Visual QA Visual QA
Languages 5 11 31
Granularity Specific indigenous Broad facets Broad categories
Parallelism Low (cultural-specific) Low (cultural-specific) Low (region-specific)
Bias Target Reasoning Failure Knowledge Gaps Knowledge Gaps
(a) Comparison of established cultural VQA benchmarks.
Feature RAVENEA [li2025ravenea] 3XCM (Ours)
Primary Task VQA & Image Cap. Cross-Modal Retrieval
Languages 8 (Countries) 16
Granularity Varies (VQA:cultural-specific, IC: too board) Universal basic-level
Parallelism Low (Mixed) High (Cross-cultural)
Bias Target Cultural Nuances Association Bias
(b) Comparison with recent multi-task and cross-modal benchmarks.
Table 4: Comprehensive Benchmark Comparison.
Appendix B Language Resources
To estimate language resource availability, we utilized the Distribution of Languages from the Common Crawl dataset (CC-MAIN-2025-18) [CommonCrawlLanguages] as an approximation. Table 5 and 6 present the resulting language composition for RQ1 and RQ2, respectively.
Type Language ID Distribution
@@ Appendix B Language Resources
Code Country Language Resources
US America English 43.950
GB Great Britain
AU Australia
DE Germany German 5.569
CN China Chinese 4.878
JP Japan Japanese 4.915
ES Spain Spanish 4.542
AR Argentina
FR France French 4.327
PT Portugal Portuguese 2.337
BR Brazil
SA Saudi Arabia Arabic 0.672
TH Thailand Thai 0.427
IN India Hindi 0.200
KE Kenya Swahili 0.010
NG Nigeria Yoruba 0.001
@@
Appendix C Evaluated Models and Details
In this work, we assess three distinct architectural paradigms for multimodal retrieval. The specific characteristics of these families are as follows:
  • Vision-Language Contrastive Models: These are foundational models trained primarily on English data. We include the original CLIP-L/14 as a powerful baseline, and CN-CLIP-L/14 to observe the effect of monolingual fine-tuning on a non-English corpus.
  • Cross-lingual Alignment Models: These models use knowledge distillation to explicitly align multilingual text encoders to a fixed, pre-trained vision space for multilingual capability. We evaluate two variants of m-CLIP, which use XLM-RoBERTa as the text encoder (XLM-R-L/14 and XLM-R-B/16plus).
Table 7 provides the precise aliases, full model names, parameter counts, and citations used throughout the paper.
Alias Used in Paper Full Model Name Parameter